if a 6 volt battery of external resistance 0.5ohms connected to 3 resistors of values 3 ohms,4 ohms,5 ohms.calculate the current in each resistor,the potential difference across the 3ohm resistor and across the parallel resistors.

1
Expert's answer
2022-03-08T16:59:21-0500

total resistance "R=0.5+3+\frac{4\times 5}{4+5}=5.722\Omega"

"I_{3\Omega}=I_{0.5\Omega}=6/5.722=1.049A (in \space series)"

"V_{0.5 \Omega}=1.049\times 0.5=0.5245V"

"V_{3\Omega}=1.049\times 3=3.147V"

"V_{4 \Omega \space and \space 5 \Omega}=IR=1.049\times \frac{4\times 5}{4+5}=2.33V (in \space parallel)"

"I_{4\Omega}=2.33/4=0.5825A"

"I_{5\Omega}=2.33/5=0.466A"